Find the generating function of recurrence relation a_(n+1) - a_n = 3n ,n<0 where ao=1


1
Expert's answer
2021-10-28T16:33:01-0400

"a_{n+1}-a_n=3n"


 characteristic equation:

"1\/x-1\/x^2=0"

"x-1=0"

"x=1"


homogeneous solution:

"a_h=cx^n=c"


particular solution:

"a_t=An^2+Bn+C"

"A(n+1)^2+B(n+1)+C-An^2-Bn-C=3n"

"2An+A+B=3n"

"A=1.5,B=-1.5"

"a_t=1.5n^2-1.5n"


"a_n=a_h+a_t=c+1.5n^2-1.5n"

"a_0=c+1.5(-1)^2-1.5(-1)=1"

"c=-2"


"a_n=-2+1.5n^2-1.5n"